OpenStudy (anonymous):

The co-ordinate of the circumcentre of triangle ABC with vertices A(3,0), B(-1,-6) and C(4, -1) is a) (1,-3) b) (-1,3) c) (3,-2) d) (-2,3)

OpenStudy (anonymous):

Reply quickly, plzzzzzzzzzz

OpenStudy (anonymous):

make the co-ordiniate of the circumcentre O (x,y)  make E(1,-3), F(1.5,-3.5) vector AB=(-4,-6) vector BC=(5,5) vector OE=(1-x,-3-y) vector OF=(1.5-x, -3.5-y) since AB and OE are perpendicular since BC and OF are perpendicular so vector AB * vector OE=0 so vector BV * vector OF =0 so (-4(1-x)+(-6)(-3-y)=0 (5(1.5-x)+(5)(-3.5-y)=0 get x and y

OpenStudy (anonymous):

x=1, y=-3

OpenStudy (anonymous):

so the coordinate of the circumcentre of triangle ABC is (1,-3)